### v1

# Detecting Neurocognitive Disorders through Analyses of Topic Evolution and Cross-modal Consistency in Visual-Stimulated Narratives

Abstract
Early detection of neurocognitive disorders (NCDs) is crucial for timely intervention and disease management. Since language abilities are affected by NCDs at a very early stage, speech analysis emerges as a non-intrusive, and scalable method for screening NCDs, especially through narrative tasks that hold significant potential for evaluating cognitive abilities. Traditional narrative analysis often focuses on local indicators in microstructure, such as word choice and syntax. While these features provide insights into language production abilities, they often fail to capture global narrative patterns, or microstructures. Macrostructures include coherence, thematic organization, and logical progressions, reflecting essential cognitive skills potentially critical for recognizing NCDs. Addressing this gap, we propose to investigate specific cognitive and linguistic challenges by analyzing topical shifts, temporal dynamics, and the coherence of narratives over time, aiming to reveal cognitive deficits by identifying narrative impairments, and exploring their impact on communication and cognition. The investigation is based on the CU-MARVEL Rabbit Story corpus, which comprises recordings of a story-telling task from 758 older adults. We developed two approaches: the Dynamic Topic Models (DTM)-based temporal analysis to examine the evolution of topics over time, and the Text-Image Temporal Alignment Network (TITAN) to evaluate the coherence between spoken narratives and visual stimuli. DTM-based approach validated the effectiveness of dynamic topic consistency as a macrostructural metric (F1=0.61, AUC=0.78). The TITAN approach achieved the highest performance (F1=0.72, AUC=0.81), surpassing established microstructural and macrostructural feature sets. Cross-comparison and regression tasks further demonstrated the effectiveness of proposed dynamic macrostructural modeling approaches for NCD detection.
Index Terms:
Neurocognitive disorder detection, Alzheimer’s Disease, macrostructure, temporal narrative analysis, sequential image storytelling, natural language processing, image processing, multi-modal recognition.
publicationid: pubid: $0$–$00$ YEAR IEEE
I Introduction
Neurocognitive disorders (NCDs)[1], also known as dementia, pose significant challenges to individuals and healthcare systems in our aging society. Early detection of these conditions is critical, as it can allow for timely intervention, potentially slowing disease progression and improving the quality of life for patients[2]. Traditional clinical approaches such as neuroimaging [3, 4], retinal photographs [5], electroencephalography (EEG) [6], cerebrospinal fluid analysis [7], and in-person neuropsychological assessments [8] have been widely investigated. However, these approaches can be resource-intensive, invasive, and costly [9, 10]. Consequently, there is a growing emphasis on developing non-invasive, cost-effective, and scalable methods for early NCD detection.
One promising approach is the automatic analysis of speech and language, as communicative impairments are often among the earliest symptoms of cognitive decline [11]. These impairments may include temporal changes in spontaneous speech [12], word finding and retrieval difficulties [13], verbal dysfluencies [14], reduction in discourse-level processing abilities [15, 16]. In light of this, various spoken cognitive tests have been designed to discover spoken language correlates for NCD detection. Such changes in speech patterns and language use provide valuable insights into underlying cognitive deficits and serve as potential biomarkers for early neurocognitive disorder (NCD) detection. In light of this, various spoken cognitive tests have been developed to uncover spoken language correlates for NCD detection. Traditional tests mostly focus on specific, discrete language functions, such as counting[17] and naming[18], where participants are required to perform highly structured and narrowly defined tasks. While these tests are useful for evaluating isolated aspects of language ability, they are limited in their scope and do not allow participants to fully demonstrate their comprehensive language capabilities. Moreover, the constrained nature of these tasks may fail to capture the dynamic, context-driven aspects of language use that are critical for assessing real-world communication skills. This highlights the need for holistic approaches to language assessment in naturalistic settings. More open-ended tasks, such as the visual-stimulated narrative (VSN) task, provide a versatile method for evaluating spontaneous language in individuals for assessment of cognitive functioning.
The VSN task involves describing a single image or telling a story based on a sequence of images, capturing an array of narrative abilities [19, 20, 21]. These narratives are deeply tied to neurocognitive domains through processes like describing, sequencing, and integrating elements. Producing a coherent narrative requires multiple cognitive functions, including attention, visual comprehension, language fluency, executive planning, memory, and social cognition [22, 23, 24, 25, 26, 27, 28]. This high cognitive demand poses challenges for individuals with NCDs, who often have deficits in these areas  [29]. Consequently, VSNs can potentially be used as a critical tool for early NCD detection [30, 11].
Approaches to narrative analysis can be categorized into two complementary aspects  [31]. First, the mainstream approach is microstructural narrative analysis, which examines linguistic (e.g., lexicon, semantics, syntax) [32, 33] and acoustic (e.g., prosody, phonology) [34, 35] features. In contrast, macrostructural analysis evaluates the overall structure and organization of storyline, emphasizing coherence (logical flow of concepts) and consistency (alignment with visual stimuli) [36]. Clinical research indicates that individuals with NCDs often exhibit greater impairment in macrostructure than microstructure, given that the microstructure provides more insights into their language abilities (perception and production), while the macrostructure offers more information on the cognitive abilities (attention, memory, organization, etc.) [31]. To date, macrostructural analysis remains underexplored, particularly within the context of NCD narratives.
Macrostructural indicators like coherence and consistency are challenging to quantify due to their abstract and conceptual nature. These indicators involve assessing how well narrative elements align with visual stimuli and how logically topics are interconnected and evolved along story progression, requiring complex cognitive processing. Traditional approaches to coherence rely on manual assessments, which are subjective and labor-intensive [37]. Recent advances in natural language processing (NLP) have introduced automated methods, such as averaging embedding similarities or using large language models (LLMs) as evaluators [38, 39]. However, these methods often fall short for NCD narratives. Embedding-based similarity can misinterpret frequent repetitions or fillers as better cohesion and ignore dynamic evolutions of narrative topics over time. LLM-generated scores lack interpretability, undermining reliability in clinical applications.
For consistency between narratives and visual stimuli, recent studies have used text-visual pretrained models to identify the highest relevance scores between image regions and sentences [40] or pattern-matching with fixed references [41, 42]. However, these methods have limitations: local text-image relevance fails to capture global consistency, and pattern-matching is restricted to fixed-reference tasks, making it unsuitable for creative tasks like storytelling. Therefore, tailored and automated approaches are needed for macrostructural analysis of NCD narratives.
This work proposed to investigate specific cognitive and linguistic challenges by analyzing topical shifts, their temporal dynamics, and the coherence of narratives over time. We presented two novel approaches to macrostructural narrative analysis: Dynamic Topic Models (DTM) for temporal analysis of topical evolution; and Text-Image Temporal Alignment Model (TITAN) for cross-modal (i.e. stimulus images and spoken narratives) alignment analysis. This approach aims to reveal cognitive deficits by identifying narrative impairments associated with NCDs, and to provide a deeper understanding of their impact on communication and cognition. The investigations are conducted using the Rabbit Story task (a VSN task) from the CU-MARVEL corpus (a Cantonese corpus), as well as the single-image description task in the ADReSS corpus (an English corpus). Our experiments validated the effectiveness of dynamic topic consistency as a macrostructure metric (F1=0.61, AUC=0.78). The proposed TITAN model achieved the highest performance (F1=0.72, AUC=0.81) in classification tasks, surpassing established microstructure and macrostructure feature sets. Moreover, experiments of cross-comparison and regression tasks further demonstrated the effectiveness of proposed dynamic macrostructure modeling approaches for NCD detection.
The paper is organized as follows: section II reviews related works, section III introduces corpora used in this study, section IV introduce the proposed approaches, section V presents our experimental settings and results, section VI analysis and discusses experimental results, and section VII concludes this work and prospects future works.
II Related Work
Extensive research has been conducted for NCD detection based on spoken language, leveraging various methodologies to identify biomarkers and assess cognitive impairments. This section reviews pertinent literature encompassing both microstructural and macrostructural narrative analysis, collectively laying the groundwork for enhanced narrative analyses. The term “microstructure” refers to fine-grained elements of narrative production, including lexical, syntax, semantics, phonetics, or acoustics, and other smaller linguistic units. In contrast, “macrostructure” refers to higher-level aspects of narrative construction, such as coherence, cohesion, thematic development, and the overall organization of narratives. Both microstructural and macrostructural analysis are crucial for a comprehensive understanding of cognitive deficits reflected in NCD narratives. However, existing studies often face limitations in capturing the full spectrum of narrative features, particularly the dynamic aspects of narrative construction. By identifying these limitations, this review highlights the gaps the proposed methods aim to fill, specifically by incorporating temporal dynamics and macrostructural consistency, thus advancing the narrative analysis of NCDs.
II-A Microstructural Analysis in NCD Detection
Individuals with NCDs exhibit speech deficits that impact spoken language across both acoustic and linguistic dimensions. Prior research has explored the various aspects of acoustic and linguistic features, encompassing prosody (e.g., dysfluencies and pauses), phonetics (e.g., Mel Frequency Cepstral Coefficients), phonology (e.g., mispronunciation of words) , voice quality (e.g., jitter and shimmer), lexicon (e.g., type-token-ratio and stopword ratio), syntax (e.g., parse-tree depth), and semantics (e.g., word and sentence embeddings) [32, 34, 43, 44, 12, 45, 46, 33, 47, 48, 41, 49, 50, 51]. These features at the microstructural level are essential for understanding the intricate language challenges faced by individuals with NCDs and serve as the foundation for microstructural analysis in this domain.
Recent efforts are devoted to identifying highly correlated and significant biomarkers of NCDs from the aforementioned features [52, 53, 54, 55]. For example, Weiner and Frankenberg et al. [53, 55] explored many traditional acoustic and linguistic features with a nested forward feature selection method, and found that the features on parts-of-speech, word categories and pauses are highly related to NCDs. While microstructural analysis provides valuable insights into specific acoustic and linguistic markers, it predominantly focuses on local features, and may overlook the broader narrative context.
II-B Macrostructural Analysis in NCD Detection
While microstructural analysis highlights the specific language production impairments associated with NCDs, macrostructural analysis is equally important as it reveals how cognitive deficits impact higher-order cognitive processes such as attention, working memory, planning, organizing and structuring, which are crucial for constructing coherent, cohesive, and thematically consistent narratives [56, 57, 58, 31].
Macrostructural analysis in the literature generally focuses on features about cohesion (the use of linguistic devices for structural unity) and coherence (logical flow and overall consistency). Individuals with NCDs tend to produce text with impaired lexical cohesion (i.e., lexical repetition or substitution), wrong co-referencing, inconsistencies in maintaining the main topic, uncompleted or irrelevant information, and disrupted sequencing of events [59, 60, 61, 62, 63].
Traditional macrostructural features were usually extracted manually based on expert experience, such as readability [64], relationship between a proposition and previous one or the global topic [65], or connection between events, macro-propositions and main information units [66].
Recent advances leverage machine learning and large language models (LLMs) for automatic macrostructure analysis on large corpora. The Coh-Metrix-Dementia program, with features like latent semantic analysis (LSA) and syntactic complexity, is well-established for extracting coherence metrics [27]. Some studies quantified global semantic coherence by using cosine similarity of word embeddings [38, 45], or evaluated pragmatic coherence via BLEU and METEOR scores using references [42]. More recent work adopted LLM as evaluator to automatically predict macro-descriptors [39].
However, as mentioned in Section I, these methods are unsuitable to NCD narratives due to frequent repetitions, unlimited references, and interpretability requirements. Additionally, there has been limited investigation into the temporal dynamics of topical evolution and the cross-modality alignment between visual stimuli and speech in VSN tasks. These gaps highlight the need for more advanced approaches to evaluate NCD narratives from both microstructural and temporal perspectives, addressing the interplay between evolving topics and the integration of visual and spoken elements.
III Corpora
This study is conducted based on the English ADReSS (Alzheimer’s Dementia Recognition through Spontaneous Speech) Corpus [67], as well as the Cantonese CU-MARVEL-RABBIT (Cognitive Assessment Using Machine LeArning EmpoweRed VoicE AnaLysis – Rabbit Story Part) Corpus. ADReSS captures speech recorded from the Cookie Theft Picture Description Task [68]. CU-MARVEL captures narratives on the Rabbit Story – a storytelling task [21].
III-A ADReSS
Figure 1: The Cookie Theft picture, where two kids are reaching for cookies, while a distracted mother is washing dishes, amidst a complex background.
The ADReSS Corpus, developed by the University of Pittsburgh, contains transcripts of spontaneous speech samples from elderly English-speaking adults. The corpus is designed to facilitate research on age-related cognitive decline, focusing on NCDs (specifically, Alzheimer’s disease and mild cognitive impairment). The neurocognitive assessment tasks in this corpus involve the Cookie-Theft Picture Description (as shown in Fig. 1) for both HCs and NCDs groups. This corpus consists of audio recordings and transcriptions of interviews with 156 participants in the Cookie-Theft Picture Description task, balanced partitioned into training (54 HCs and 54 NCDs) and testing (24 HCs and 24 NCDs) subsets.
III-B CU-MARVEL’s Rabbit Story Task
The CU-MARVEL Corpus is designed and collected by our group in Hong Kong, with the participant inclusion criteria: (i) aged 60 or above; (ii) sufficient Cantonese language skills (listening and speaking), and (iii) sufficient vision and hearing ability (where glasses and hearing aids are acceptable) to complete all cognitive tests. This corpus contains series of cognitive tests, to align with the picture description task, we adopt the speech samples in the Rabbit Story task of CU-MARVEL, denoted as “CU-MARVEL-RABBIT” Corpus. It records Cantonese-speaking participants narrating a story about a girl finding a rabbit through 15 cartoon images, as illustrated in Fig. 2. These images are presented in a booklet, with each image displayed on a separate page. Participants will begin by reviewing all 15 images to familiarize themselves with the content. Subsequently, they will examine each image individually, providing a description of its content before proceeding to the next image. This process will continue sequentially until descriptions for all 15 images have been completed.
Figure 2: The Rabbit Story cartoon images – showing how a girl and her dog were trying to find a rabbit, and encountering various animals and adventures throughout their journey.
The CU-MARVEL-RABBIT Corpus differs from the ADReSS Corpus in its task design and analytical focus. While the ADReSS Corpus centers on single-image description to evaluate linguistic features associated with NCD, CU-MARVEL-Rabbit employs a multi-images description task, emphasizing the cognitive ability to connect sequential images and narrate a logical story. This task prioritizes macrostructural analysis, including coherence and cohesion, to assess narrative organization at a higher level. The version “2024-06-13” of the CU-MARVEL-RABBIT Corpus consists of 758 subjects (467 HCs, 291 NCDs), where the data distribution and the train-test split are listed in the Table I. The CU-MARVEL-RABBIT Corpus records the assessment for each subject based on specialists in neurology, along a scale from 0-4 regarding cognitive functioning, where the values 0 and 1 indicate HCs and 2, 3, and 4 indicate Neurocognitive Disorders (NCDs).
TABLE I: Distribution of CU-MARVEL-Rabbit Corpus
Split Gender Label Total
0 1 2 3 4
Train Male 94 107 5 86 12 304
Female 81 145 37 74 17 354
Total 175 252 42 160 29 658
Test Male 10 10 10 10 10 50
Female 10 10 10 10 10 50
Total 20 20 20 20 20 100
IV Approach
In this work, we propose a comprehensive analysis for the automatic detection of Neurocognitive Disorders based on visual-stimulated narratives, encompassing both microstructural and macrostructural analysis. Microstructural analysis, detailed in Section IV-A, focuses on the extraction of acoustic and linguistic features selected from previous literature. On the other hand, macrostructural analysis, presented in Sections IV-B to IV-D, introduces three distinct methodologies to capture global and temporal patterns with varying focal points.
TABLE II: Microstructural features used in this work.
“%”: “ratio of”, “#”: “number of”, “dur”: “duration”.
ID Feature Name Description
Acoustics
A1 # Total pause # pauses
A2 Total pause dur all pause dur in seconds
A3 Average pause dur averaged pause dur in seconds
A4 Normalized pause dur Pause dur / articulation rate
A5 Pause frequency # pauses / speech dur
A6 Pause occurrence rate # pauses / # characters
A7 Total utterance dur all utterance dur in seconds
A8 Average utterance dur averaged utterance dur in seconds
A9 Articulation rate # characters spoken in a second
Linguistics
L1 % Stop words # stop words / # words
L2 % Filler pauses # filler words / # words
L3 % Lexical filler # lexical filler / # words
L4 % Repetition # repeated words / # words
L5 % Pronoun verb change # pronoun to verb changes / # words
L6 % Pronoun noun change # pronoun to noun changes / # words
L7 % Pronoun # pronouns / # words
L8 % Verb # verb / # words
L9 % Noun # nouns / # words
L10 % Adj # adjectives / # words
L11 % Pronoun verb # pronouns / # verbs
L12 % Pronoun noun # pronouns / # noun
L13 % Type of token (TTR) # different tokens / # tokens
IV-A Traditional Acoustic and Linguistic Analysis
Based on the literature review, the microstructural features have been categorized into two main groups: acoustic features and linguistic features. We select 9 acoustic features and 13 linguistic features from the literature that are suitable for Cantonese and CU-MARVEL-RABBIT Corpus, as summarized in Table II. Before extracting these features, we preprocess all audios with FullSubNet+ [69] denoiser and peak-normalization to ensure internal consistency and avoid Clever Hans Effect [70].
The rows of A1-A9 in Table II show the name and the description of the acoustic features, related to pause (A1-A6), and speech duration (A7-A8) and speaking rate (A9). To extract these features, we first employ Voice Activity Detection (VAD) using the Feedforward Sequential Memory Network, as described in the paper [71], to obtain the timestamps of the voiced segments. With these segment timestamps, a pause is detected by the silence gap between two adjacent voiced segments when the duration exceeds the duration of a syllable for this speaker, which is calculated by dividing the total voiced duration by the total number of characters for each Cantonese participant.
The linguistic features are extracted from the automatic transcriptions of participants obtained using the ASR model SenseVoice-Small [72]. The rows of L1-L13 in Table II show the name and the corresponding description of each feature. To extract these linguistic features from Cantonese transcripts, we utilized PyCantonese [^1] https://pycantonese.org to do word segmentation and POS auto-tagging, combining vocabulary from Jieba [^2] https://github.com/fxsjy/jieba.
Figure 3: The topic distribution for the HC group (a) and the NCD group (b), and the topic evolution for the HC group (c) and the NCD group (d), as obtained from DTM. (a-b) show the weights of each topic word (top-10) within the topic (from A to E) for the group. (c-d) display average topic proportions with curves and deviations with filled areas. The order of topics A to E is based on the positions of their first peaks in topic evolutions.
IV-B Reference-based Narrative Alignment
A critical part of evaluating the narrative task, particularly within the Rabbit Story context, involves assessing the comprehensiveness of a subject’s narrative. To this end, we identify keywords associated with the task and measure the coverage in the subject’s narratives. Specifically, we leverage the TF-IDF approach from information retrieval to compute the relative importance of each word within HC groups. After selecting the top 100 words based on TF-IDF weights and excluding punctuation and stopwords, these words are categorized into four semantic groups informed by human prior knowledge of the Rabbit Story and the Narrative Scoring Scheme (NSS) proposed by [24]. The groups include introduction setting words (e.g., one day, morning), character words (e.g., girl, dog, rabbit), object words (e.g., park, home, stone), and action words (e.g., play, find, walk). Coverage (abbreviated as “cvg”) for each category is then computed as features for analysis. While these coverage metrics effectively capture whether local narrative elements are addressed, they lack insights into global coherence and connections across the entire picture set. To address this limitation, metrics commonly used in image captioning tasks, such as BLEU-n (n=1,2,3,4), METEOR, and ROUGE-L, are employed to quantify the degree of alignment between subject narratives and reference descriptions. For this purpose, we engaged six colleagues to craft curated reference descriptions of the Rabbit Story that align with the rubric of NSS, in the aspects of introduction, character development, mental and emotion states, referencing, event&reaction, cohesion, and conclusion. The final BLEU-n, METEOR, and ROUGE-L scores are averaged across these references for each subject, providing additional features that reflect global narrative coherence. With these reference-based features, we can perform macrostructural analysis that evaluates the coherence against some gold standards; however, these features do not account for temporal changes in topics or the cohesion of narrative over time, which is another important part of microstructural analysis.
IV-C DTM-based Temporal Analysis
Temporal changes or thematic evolution in narratives are critical for understanding the progression and manifestations of NCDs, as these elements reflect higher-level cognitive abilities. To capture these temporal dynamics, we utilize Dynamic Topic Modeling (DTM) [73], which tracks them by explicitly modeling the evolution of topics over time. In contrast to previously mentioned static methods, DTM offers a dynamic perspective on the underlying temporal trends within the narratives.
To understand how DTM captures these temporal dynamics, it is important to first grasp the basics of topic modeling. Topic model provides a probabilistic framework to undercover the latent semantic structure within textural data. While the arbitrarily generated text may randomly be distributed across documents, the topic model assumes that words in natural languages are distributed around meaningful semantic topics. Specifically, document generation is assumed to be based on a probabilistic combination of underlying topics, where each topic is represented by a probabilistic distribution over vocabulary words. In the classical Latent Dirichlet Allocation (LDA), these distributions are assumed to be static and drawn from Dirichlet priors.
Dynamic Topic Models (DTM) is an extension of classical topic models, it introduces time dependencies upon the distributions, making it particularly suitable for analyzing temporal changes in narrative content. Formally, DTM models the dynamic nature of topic evolution by introducing time dependency $t$ in topic word distribution $\beta_{k,t}$, where $k$ and $t$ denote the topic and time indices respectively. In this way, it allows the topic distributions to drift over time via a Gaussian process, i.e., $\beta_{t}|\beta_{t-1}\sim\mathcal{N}(\beta_{t-1},\boldsymbol{\Sigma})$, where $\boldsymbol{\Sigma}$ represents the covariance matrix controlling topic transition smoothness.
Then the generative process of DTM proceeds as follows:
Firstly, for each document $d$ in time slice $t$, the model first samples a topic proportion from a Dirichlet prior: $\theta_{d}\sim\text{Dir}(\alpha)$, where $\alpha$ is the Dirichlet concentration parameter controlling the sparsity of topic distributions.
Secondly, for each word position $n$ in document $d$, the model samples a topic assignment from a multinomial distribution parameterized by the document topic proportion: $z_{d,n}\sim\text{Mult}(\theta_{d})$.
Finally, the word is sampled from a multinomial distribution defined by the assigned topic’s word distribution at that specific time slice: $w_{d,n}\sim\text{Mult}(\beta_{t,z_{d,n}})$.
The model parameters could be estimated through variational inference [74] or the Expectation-Maximization [75] algorithms during the training phase.
In order to apply the DTM model to narrative analysis for NCD detection, we evenly split each participant’s ASR transcription into 15 slices, assuming that each image leads to approximately similar words of spoken narratives. This segmentation allows us to track the evolution of participants’ narrative topics across the story’s progression, enabling a granular analysis of discourse cohesion by observing how thematic content shifts and develops with each sequential image in the narrative.
We use DTM by inputting the segmented transcription data, where each slice represents a distinct portion of the narrative associated with a time step. The DTM processes this data to identify evolving topics within and across the segmented slices, allowing us to observe how smoothly and coherently topics transition as the story progresses.
Fig. 3 illustrates the topic distribution (a, b) and temporal evolution of topic proportions (c, d) for the HC and NCD groups. For the distribution of topics over words, there exists differences between HC and NCD groups, where HC topics over time tend to be more aligned with the order of the images. For example, the topic words of the HC group, shown in Fig. 3(a), include “house” and “garden” in topic “A”, “ant” and “tree” in topic “B”, “goose” and “caw” in topic “C”, “pond” in topic “D”, and “many” “rabbits” in topic “E”. These topic words generally align with the different scenes along the story. In contrast, the topic words of the NCD group, shown in Fig. 3(b), are more chaotic, e.g., topic “A” and “B” are mostly overlapped with “rabbit” and “dog”, while “stable” in “C” and “fall” in “E” are both derived from images 10-11.
Another interesting finding is that the HC narratives exhibit a cyclical pattern. Specifically, most HC individuals mention “return” and “home” in topic ”E”, even though these elements are not depicted in the final images. This implies that HCs can tie the ending of the story back to its beginning, completing the girl’s quest to find the lost rabbit and bringing the story to a satisfying closure. For the dynamic proportions of these topics over time, i.e., the Gaussian Process, it is observable that the HC group exhibits more distinct peaks and smoother transitions between topics over time, while the NCD group shows more erratic changes with less pronounced peaks.
Based on these findings, we proposed several DTM-related statistics for NCD detection. The first metric, termed “DTM-based topic consistency”, is derived from the hit rate of the top-10 topic words each time, which are derived from a DTM pre-trained on all narratives. This statistic measures how consistently the subject’s topic words align with the top-10 words over time, thus providing insights into the stability and coherence of topics across different groups (HC vs. NCD). Higher topic consistency suggests a more stable thematic structure, indicating a potentially healthier cognitive state. The second metric, termed “DTM-based cyclical pattern”, which is simply a boolean value indicating whether the word “home” appears in the latter half of each narrative.
We also proposed several metrics by calculating the dynamic properties of topic evolution, including standard deviation of the proportions over time, peak-to-peak range (distance between positions of the highest and lowest points), and rate of change (averaged Euclidean distance of adjacent topic distributions over time).
IV-D TITAN: Text-Image Temporal Alignment Network
While DTM demonstrates potential in capturing topic evolution patterns, it faces challenges in incorporating cross-modal information in tasks such as picture description and visual-stimulated narratives. This limitation impairs topic representation model, preventing the effective utilization of image information and the handling of cross-modality correspondence. Previous studies have attempted to address these issues by inputting images into multimodal large language models to generate reference descriptions and keywords [42]. However, these approaches often overlook temporal dynamics and fail to capture topic evolution throughout the entire narrative. To overcome these challenges, and drawing inspiration from multimodal metrics like CLIPScore and GROOViST [76, 77], we propose the Text-Image Temporal Alignment Network (TITAN). TITAN treats images as dynamic topics and constructs temporal relationships using the Attention mechanism, as illustrated in Fig. 4.
This network comprises several key components:
D1. Feature Extraction
The first step involves extracting features from the input text and images using the pretrained text-visual models, comprising text-tower and visual-tower to process different modalities. In this work, we adopted the Chinese-CLIP [78] model, which consists of visual-tower ViT and text-tower RoBERTa. Chinese-CLIP is pre-trained in a contrastive learning paradigm that aims to match the images with their corresponding descriptions in a multi-modal embedding space, specifically tailored for the Chinese language (including Cantonese). We adopt this model because it can effectively align the embeddings of images and Cantonese text in the semantic space, thus providing insights for narrative coherence (structural dynamics) and thematic consistency (content alignment). Specifically, given an input image $I$ and text $T$, we obtain the respective embeddings $E_{I}$ and $E_{T}$ as follows:
$E_{I}=\text{Chinese-CLIP}_{ViT}(I)\in\mathbb{R}^{J\times H}$
$E_{T}=\text{Chinese-CLIP}_{RoBERTa}(T)\in\mathbb{R}^{K\times H}$
where $H$ is the hidden dimension of the extracted embeddings, $K$ is the temporal length of the investigated text sequence, and $J$ is the number of image frames, e.g., $J=1$ for the ADReSS dataset and $J=15$ for the CU-MARVEL-RABBIT corpus.
Figure 4: (a) The proposed model TITAN. (b) The Attention module of TITAN. (c) The correlation map between image and text embeddings in TITAN. TITAN takes images and text narratives as inputs to predict NCDs, with Attention using RoPE to focus on useful positions.
D2. Embedding Fusion
The extracted embeddings $E_{I}$ and $E_{T}$ are then concatenated to form a combined representation:
$E_{IT}=[E_{I};E_{T}]\in\mathbb{R}^{(J+K)\times H}$
where [;][;][ ; ] denotes the concatenation operation. This concatenation operation offers at least two advantages. First, it allows the subsequent layers to share the same projection layers for both image and textual embeddings, which is parameter-efficient and feasible since the embeddings are aligned to a common multimodal space during the pre-training stage. Second, inputting the concatenated embeddings into the Attention mechanism enables the model to capture combined information from image-image, image-text, and text-text pairs, thereby enhancing the richness of the contextual features.
D3. Projection Layers
The combined embedding $E_{IT}$ is dimensionally adjusted through multiple projection layers, starting with a linear projection to reduce the embedding size:
$E_{down}=E_{IT}W_{down}^{T}+b_{down}\in\mathbb{R}^{(J+K)\times H^{\prime}}$
where $H^{\prime}$ is the reduced dimension, $W_{down}\in\mathbb{R}^{H^{\prime}\times H}$ and $b_{down}\in\mathbb{R}^{H^{\prime}}$ are the weight matrix and bias of the projection layer, respectively.
After the Attention module, the embedding is then up-projected to its original dimension for residual connections:
$E_{up}=E_{attn}W_{up}^{T}+b_{up}\in\mathbb{R}^{(J+K)\times H}$
where $W_{up}\in\mathbb{R}^{H\times H^{\prime}}$ and $b_{up}\in\mathbb{R}^{H}$ are the parameters of the up-projection layer.
This bottleneck projection structure effectively balances dimensional reduction for computational efficiency and dimensional expansion for capturing complex feature interactions.
D4. Attention Module
As shown in Fig. 4(b), the Attention module aims to capture the temporal dynamics and contextual relationships within and between modalities, by the attending on specific positions of the projected embeddings.
To effectively process dynamic temporal information, we incorporate Rotary Position Embedding (RoPE) [79] with the Attention mechanism. RoPE encodes positional information into token embeddings without disrupting the invariance properties of self-attention, making it well-suited for capturing temporal dependencies. Mathematically, the positional encoding vector for the $i$-th position, $p_{i}\in\mathbb{R}^{2H^{\prime}}$, is defined using sinusoidal functions:
$p_{i}=[\sin(i),\cos(i),\ldots,\sin(\frac{2H^{\prime}}{2}i),\cos(\frac{2H^{\prime}}{2}i)]^{T}$
Each element of $p_{i}$ is derived from sinusoidal functions to capture the position information: $\sin(i)$ and $\cos(i)$ are applied to alternate dimensions, with increasing frequencies across dimensions to capture information at various scales. Given the different temporal strides of images and texts, RoPE is applied separately to each modality with different frequencies, i.e., $p_{i}\in\mathbb{R}^{1,\ldots,H^{\prime}}$ for images and $p_{i}\in\mathbb{R}^{H^{\prime}+1,\ldots,2H^{\prime}}$ for texts.
Given the projected embedding $E_{down}$, we compute the queries, keys, and values as follows:
$Q=E_{down}W_{Q}^{T},K=E_{down}W_{K}^{T},V=E_{down}W_{V}^{T}$
where $W_{Q},W_{K},W_{V}\in\mathbb{R}^{H^{\prime}\times H^{\prime}}$ are learnable parameter matrices for query, key, and value projections respectively. The queries and keys, for image and text parts separately, are then transformed using RoPE to compute position-aware Attention scores. Specifically, for each query $q_{i}$ and key $k_{i}$, the transformed versions are computed as:
$q_{i}^{\prime}=q_{i}\cos(p_{i})+q_{i}^{\perp}\sin(p_{i}),k_{i}^{\prime}=k_{i}\cos(p_{i})+k_{i}^{\perp}\sin(p_{i})$
where $q_{i}^{\perp}$ and $k_{i}^{\perp}$ represent the perpendicular components of the query and key vectors.
Finally, the dot-product attention is computed as follows:
$E_{attn}=\text{softmax}\left(\frac{\text{Mask}(Q^{\prime}K^{\prime T})}{\sqrt{H^{\prime}}}\right)V$
where $Q^{\prime}$ and $K^{\prime}$ are the RoPE-transformed queries and keys. Mask ensures the padding positions of input get near-zero attention weights after the softmax. The scaling factor $\sqrt{H^{\prime}}$ prevents the dot products from growing too large in magnitude. The values remain untransformed to retain the original contents that are being passed through the network.
In visual-stimulated narrative tasks such as Rabbit Story, narrative transcriptions must align with reference images. Participants with NCDs may misidentify animal names (anomia) or miss detailed scenarios (visual perception) in these images. For example, most NCDs misidentify “goose” as “duck”, or skip the event of “goose in the brush” in the seventh image. The attention module in proposed text-image temporal alignment network is able to perceive potential misalignment and discrepancies cross modalities, and therefore recognize NCD participants with the consideration of their narrative macrostructure.
TABLE III: Classification and Regression Results on CU-MARVEL-RABBIT. The quantity or dimension of the features is shown in parentheses. The star ‘*’ in ‘Sys.’ signifies the baseline systems. $\downarrow$ indicates lower is better. “Emb.” denotes embedding(s).
Sys. Features Model Classification Regression
[…]
1* Acoustics (9) SVM 0.5778 0.8029 0.4333 0.8667 0.6200 0.178 0.3206
2* Linguistics (13) SVM 0.2933 0.7296 0.1833 0.7333 0.4700 -0.0377 0.3602
3* Micro (22) SVM 0.5843 0.8204 0.4333 0.8966 0.6300 0.1904 0.3181
4 Reference-based (16) SVM 0.6735 0.7350 0.5500 0.8684 0.6800 0.2381 0.3086
5 DTM-based (5) SVM 0.6737 0.7792 0.5333 0.9143 0.6900 0.2508 0.3060
6 Macro (21) SVM 0.6804 0.7821 0.5500 0.8919 0.6900 0.2864 0.2987
7 All statistics (43) SVM 0.6947 0.7942 0.5500 0.9429 0.7100 0.2570 0.3048
8 Text-Visual Emb. (1024) TITAN 0.7238 0.8120 0.6333 0.8444 0.7100 0.3876 0.3130
D5. Prediction
Finally, the up-projected embedding is temporal pooled with Attention mask ($\in\mathbb{R}^{J+K}$) and passed through a fully connected layer to generate the prediction:
$E_{pool}=\frac{\sum_{i=1}^{J+K}\text{mask}_{i}\cdot E_{up,i}}{\sum_{i=1}^{J+K}\text{mask}_{i}}$
$\hat{y}=\sigma(E_{pool}W_{fc}^{T}+b_{fc})\in\mathbb{R}^{C}$
where $C$ is the prediction dimension (number of classes for classification tasks and 1 for the regression task), $W_{fc}\in\mathbb{R}^{C\times H}$ and $b_{fc}\in\mathbb{R}^{C}$ are the fully connected layer’s weight and bias, and $\sigma$ is an activation function (softmax for classification tasks and sigmoid for the regression task).
Throughout the network, residual connections are employed at each stage to facilitate gradient flow and mitigate the vanishing gradient problem.
The TITAN model, with its robust mechanisms for feature extraction, embedding fusion, and Attention-based contextual alignment, effectively integrates text and image information to evaluate the consistency and temporal flow in narratives. As shown in Fig. 4(c), the heatmap illustrates how the model captures the alignment between text chunks (x-axis) and a sequence of images (y-axis). The varying intensity of the colors represents the degree of alignment between the image and textual content, and the diagonal intensity represents the temporal consistency and flow of the narrative. Bright spots indicate high alignment, showing that text chunks and images correlate strongly at those points, while the diagonal progression reflects the sequence in which events or scenes unfold coherently over time. Moreover, compared to Fig. 3(c, d) using term-based topic representation, Fig. 4(c) demonstrates a more fine-grained temporal alignment between narratives and topics/images, which highlights the efficiency of representing topics with images.
V Experiments & Results
V-A Experimental Setup
In this study, we conduct both classification and regression tasks to evaluate the performance of various feature sets and classifier models. Eight systems are designed, each utilizing distinct feature sets. System 1 employs nine acoustic features, while System 2 incorporates 13 linguistic features both introduced in Section IV-A. System 4 utilizes features derived from reference-based analysis as described in Section IV-B, and System 5 employs statistical features extracted from the DTM model outlined in Section IV-C. Systems 3, 6, and 7 leverage different combinations of the feature sets used in the aforementioned systems. As these features are all statistical in nature, we first apply Principal Component Analysis (PCA) with 5 components to reduce dimensionality and then utilize a Support Vector Machine (SVM) for prediction, with a Radial Basis Function (RBF) kernel and soft margin $C=1$. The hyper-parameters “components” of PCA, and “kernel” and “soft margin” of SVM are grid searched using System 7 with all statistical features. System 8, the TITAN model, as depicted in Fig. 4, is connected with a fully-connected linear layer as the task head. To align with PCA in Systems 1 to 7, the projection dimension $H^{\prime}$ of TITAN is also set as 5.
V-B Evaluation Protocol
For the evaluation, we employed multiple metrics across both classification and regression experiments. For the classification task, we used the F1-score and AUC as the primary metrics, supplemented by recall (Rec.), precision (Prec.), and accuracy (Acc.), reported on the positive class (NCD). Despite of AUC, the classification metrics were computed with a default threshold set at 0.5. For regression, we used the coefficient of determination ($R^{2}$, ranging from $\infty$- ∞ to 1) and Root Mean Square Error (RMSE, ranging from 0 to +$\infty$+ ∞) as metrics. To ensure robust evaluation, we averaged the results from the final five epochs of each experiment. Notably, both the ADReSS and CU-MARVEL-RABBIT test sets were unseen and maintained an equal distribution of positive and negative samples, ensuring a balanced evaluation.
V-C Results
Table III presents the results of classification and regression tasks, comparing various feature sets and models using F1 as the primary evaluation metric. Among the systems, the proposed TITAN model (Sys. 8), which utilizes combined text and image embeddings, achieves the best performance in both tasks generally, with an F1 score of 0.7238 and AUC of 0.8120 in classification, and R2 of 0.3876 in the regression task. In contrast, linguistic features-based system (Sys. 2) shows an inferior performance among these systems, which is limited by the ASR, segmentation and tagger systems that are not tailored for colloquial Cantonese narratives.
Another finding is that macrostructural statistics (Sys. 6) outperform microstructural statistics (Sys. 3), achieving F1 scores of 0.6804 versus 0.5843 in classification, and R2 of 0.2864 versus 0.1904, indicating the efficacy of discourse-level analysis on NCD narratives.
VI Analysis & Discussion
Figure 5: Correlation coefficients between proposed features and NCD labels, highlighting the statistical significance of each feature on the x-axis.
VI-A Comparison with previous literature
To evaluate the performance of the models, we conducted a cross-comparison study on the semi-public ADReSS dataset. The results in Table IV demonstrates that the proposed TITAN model (Sys.8) achieves superior performance with the F1 score of 0.8889, exceeding others by at least 3.13%.
TABLE IV: Cross comparison study (binary classification) on ADReSS.
Reference F1 AUC Rec. Prec. Acc.
Yuan’2020 [80] 0.8619 - 0.8540 0.8700 0.8540
Zhu’2023 [40] - - - - 0.8224
Botelho’2024 [39] $0$ 0.8330 0.7690 0.8130
Li’2024 [42] 0.8510 - 0.8330 0.8700 0.8540
Ours (Sys.8) 0.8889 0.9410 0.8333 0.9524 0.8958
VI-B Statistical Features Correlation Ranking
To evaluate the significance of the proposed statistical features, we performed Spearman’s rank correlation test to rank all features based on the absolute values of correlation coefficients. As illustrated in Fig. 5, the correlation distribution indicates that the microstructural features demonstrate greater importance than macrostructural features. Notably, nine of the top ten features are derived from microstructural analysis, highlighting the effectiveness of these features. Specifically, DTM-based metrics such as “topic consistency,” “rate of change,” and several reference-based coverage metrics ranked relatively high with statistical significance ($p<0.001$), suggesting their potential as indicative features. Furthermore, many features associated with narrative macrostructural abilities display highly negative correlations with NCDs, whereas most pause-related features exhibit positive correlations with NCDs. These findings underscore the potential utility of these features for effective classification and a deeper understanding of NCD-related characteristics.
VI-C Text-Visual Alignment Differences between NCDs and HCs
Figure 6: Averaged correlation maps between image embeddings (y-axis) and text embeddings (x-axis) from HCs (a) and NCDs (b) using TITAN.
Fig. 6 quantifies the cross-modal alignments among individuals of HCs and NCDs. The highlighted cells in the figure represent regions with high text-visual correlation, indicating stronger alignment between the image embeddings (y-axis) and the text embeddings (x-axis). Notably, The HC narratives (Fig. 6a) exhibited consistently relatively high correlation scores across the heatmap diagonal, signifying cross-modal alignment throughout the narrative sequence. In contrast, the NCD narratives (Fig. 6b) displayed dispersed attention scores, reflecting disjointed narrative elements corresponding to the image stimuli. This distinction underscores TITAN’s ability of modeling cross-modal consistency for narrative analysis, as well as its potential in identifying and assessing cognitive decline.
VI-D Exemplar Study on NCD Detection Task
Figure 7: Example of an NCD narrative, where the Cantonese words are transcribed by ASR and English words are translated by GPT-4o.
To elaborate the investigated patterns in this work, we select a specific example with NCD from the unseen test set, as shown in Fig. 7. The Cantonese transcription is obtained using SenseVoice-Small, and the translation is provided by GPT-4o [81] to convey key meanings of it, which omits repetitions, dysfluencies and stop words. It is observable that this narrative misses many animals and events from the stimuli, such as squirrel, goose, and their related events. Additionally, there are few transitions between sentences or events during the narrating process.
To better understand these patterns, especially the proposed features, we quantify their contributions of classifying this example using the SHAP (SHapley Additive exPlanations) value [82], as shown in Fig. 8. SHAP values indicate contribution of each feature during the process of classification, showing how each feature influences the final decision starting from the classification boundary. Fig. 8 compares the contributions in the systems using microstructural and macrostructural features respectively. For the classifier using microstructural features, the top contributing feature “speech duration” tends to correctly predict the case, but most other features hinder this prediction. Conversely, the classifier using macrostructural features highlights the rate of change (topic evolution) and reference-based metrics as key indicators, which help make a correct prediction. As a result, the microstructural analysis fails to accurately detect the NCDs, whereas the macrostructural analysis successfully identifies them. These distinctions underscore the importance of macrostructural patterns in distinguishing NCDs.
VI-E Ablation Study of TITAN
We conducted an ablation study to assess the contributions of different components of the proposed TITAN model, as detailed in Table V. By removing specific parts (denoted as ‘w/o’), we observed their impact on performance for the NCD detection task. The results show that each component is essential, with the absence of the image embedding (w/o Image emb) causing the most significant performance drop, highlighting its critical role in the model’s efficacy.
TABLE V: Ablation study of TITAN (Classification)
Approach F1 AUC Rec. Prec. Acc.
TITAN (Sys. 8) 0.7238 0.8120 0.6333 0.8444 0.7100
w/o RoPE 0.7171 0.8203 0.6167 0.8566 0.7080
w/o Attention 0.6796 0.7962 0.5833 0.8140 0.6700
w/o Image emb 0.6250 0.7855 0.5000 0.8333 0.6400
(a) Microstructural (Ground-truth: 1, Prediction Probability: 0.3159)
(b) Macrostructural (Ground-truth: 1, Prediction Probability: 0.7059)
Figure 8: Feature contributions for the classifiers using (a) microstructural features and (b) macrostructural features. The x-axis and bars represent the magnitude (SHAP values) and direction (colors) of feature contributions. The y-axis lists the names and standardized values of the top features.
VII Conclusion & Future Work
Macrostructural narrative analysis, which measures global coherence and cohesion at the narrative level, is a vital research scope for NCD detection. However, existing studies typically ignore its temporal aspects and the dynamic topic evolution in NCDs narratives. In this study, we proposed two dynamic macrostructural modeling approaches, DTM-based temporal analysis and the Text-Image Temporal Alignment Network (TITAN), to measure the temporal alignment between the narrative topics and the implicit topics in references or visual stimuli. The DTM-based temporal analysis revealed the erratic topic evolution of NCDs, while TITAN further extracted the referenced dynamic topics from the sequential images in a more fine-grained manner. These approaches were evaluated on the Cantonese CU-MARVEL-RABBIT corpus for classification and regression tasks. TITAN achieved the highest performances in both tasks, surpassing the baseline feature sets and DTM-based metrics. Additional cross-comparison on the English ADReSS corpus and regression task further validated the effectiveness of the dynamic macrostructural modeling approaches.
We anticipate the temporal and macrostructural approaches to significantly contribute to the early screening and detection of NCDs, providing a more comprehensive framework for analyzing narratives from these perspectives. While these approaches can still be technically optimized, the concepts underlying them offer new insights into quantifying language use that may indicate cognitive decline, and enhance the understanding of NCDs in visual-stimulated narratives. Moreover, they can be extended to enhance multimodal fusion of more modalities, support multilingual generation due to their language-agnostic nature, and detect other spoken disorders.