### v2

# Detecting Neurocognitive Disorders through Analyses of Topic Evolution and Cross-Modal Consistency in Visual-Stimulated Narratives

Abstract
Early detection of neurocognitive disorders (NCDs) is crucial for timely intervention and disease management. Given that language impairments manifest early in NCD progression, visual-stimulated narrative (VSN)-based analysis offers a promising avenue for NCD detection. Current VSN-based NCD detection methods primarily focus on linguistic microstructures (e.g., pauses, lexical diversity), which are potentially linked to bottom-up (stimulus-driven) cognitive processing. While these features illuminate basic language abilities, the higher-order linguistic macrostructures (e.g., thematic or logical development), which may reflect top-down (concept-driven) cognitive abilities, remain underexplored. These patterns are crucial for NCD detection yet challenging to quantify due to their abstract and complex nature. To bridge this gap, we propose two novel dynamic macrostructural approaches: (1) Dynamic Topic Model (DTM) to track topic evolution over time, and (2) Text-Image Temporal Alignment Network (TITAN) to measure cross-modal consistency between speech and visual stimuli. Experimental results validated the efficiency of proposed approaches in NCD detection, with TITAN achieving superior performance both on the CU-MARVEL-RABBIT corpus (F1 = 0.7238) and the ADReSS corpus (F1 = 0.8889). The feature contribution analysis revealed that macrostructural features (e.g., topic variability, topic change rate, and topic consistency) constituted the most significant contributors in the model’s decision pathways, outperforming investigated microstructural features. These findings underscore the critical role of macrostructural patterns in understanding cognitive impairment mechanisms in NCDs.
Index Terms:
Neurocognitive disorder detection, Alzheimer’s Disease, macrostructure, temporal narrative analysis, sequential image storytelling, natural language processing, image processing, multi-modal recognition.
publicationid: pubid: $0$–$00$ YEAR IEEE
I Introduction
Neurocognitive disorders (NCDs)[1, 2], such as Alzheimer’s Disease, are highly prevalent and pose significant challenges to individuals and healthcare systems in our aging society. Early detection of these conditions is critical, as it can allow for timely intervention, potentially slowing disease progression and improving the quality of life for patients[3]. Traditional clinical approaches such as neuroimaging [4, 5], retinal photographs [6], electroencephalography (EEG) [7], cerebrospinal fluid (CSF) analysis [8], and in-person neuropsychological assessments [9] have been widely investigated. However, these approaches can be resource-intensive, invasive, and costly [10, 11]. Consequently, there is a growing emphasis on developing non-invasive, cost-effective, and scalable methods for early NCD detection.
One promising approach is the automatic analysis of speech and language, as communicative impairments are often among the earliest symptoms of cognitive decline [12]. These impairments may include temporal changes in spontaneous speech [13], word finding and retrieval difficulties [14], verbal dysfluencies [15], reduction in structural and thematic organization abilities [16, 17]. Such changes in speech patterns and language use provide valuable insights into underlying cognitive deficits and serve as potential biomarkers for early NCD detection. In light of this, various spoken cognitive tests, such as counting [18] and naming [19], have been designed to discover spoken language correlates for NCD detection. Among these, the visual-stimulated narrative (VSN) task [20, 21], such as describing a single image or telling a story based on image sequences, stands out by employing visual stimuli to elicit rich and spontaneous speech within a naturalistic and holistic framework. Specifically, the VSN task engages processes like viewing, describing, sequencing, and integrating elements, requiring multiple cognitive functions, such as attention, visual comprehension, language fluency, executive planning, memory, and social cognition [22, 23, 24, 25, 26, 27, 28]. Therefore, it offers superior potential for comprehensively evaluating cognitive abilities for NCD detection [29, 12].
Approaches to the analysis of VSNs can be categorized into two complementary aspects  [30]. First, the mainstream approach is microstructural analysis, which examines detailed, localized elements of a narrative, such as linguistic (e.g., lexicon, semantics, syntax) [31, 32] and acoustic (e.g., prosody, phonology) [33, 34] units. This level focuses on how things are produced and primarily reflects bottom-up (exogenous, stimulus-driven) cognitive processing. In contrast, macrostructural analysis evaluates the overarching organization of the narrative units, such as informativeness (overall information accuracy and completeness), coherence (logical flow of events or concepts) and text-visual consistency (alignment with visual stimuli [^1] This includes both depiction per image and connection across images, bridging informativeness and coherence.[35]. This level focuses on what is said and how things are combined, generally representing top-down (endogenous, concept-driven) cognitive integration.
Clinical research indicates that individuals with NCDs often exhibit greater impairment in macrostructuring than microstructuring. While microstructural deficits typically reveal specific language abilities (perception and production), macrostructural deficits more directly reflect impairments in higher-order cognitive abilities (attention, memory, organization, executive function), manifesting as disrupted topic maintenance, incoherent sequencing, and inconsistency to visual stimuli [30]. Despite this, macrostructural analysis remains underexplored, particularly within the context of NCD narratives.
This research gap primarily arises from the inherent difficulty in quantifying macrostructural indicators due to their abstract and conceptual nature. Traditional approaches to quantify coherence rely on manual assessments, which are subjective and labor-intensive [36]. Recent advances in natural language processing (NLP) have introduced automated methods, such as calculating cosine similarities between consecutive textual embeddings or using large language models (LLMs) as evaluators [37, 38]. However, these methods often fall short for NCD narratives. Embedding-based similarities can misinterpret frequent repetitions or fillers as better cohesion and ignore dynamic evolutions of narrative topics over time. LLM-generated scores lack interpretability due to their untraceable decision pathways, undermining reliability in clinical applications.
To quantify the consistency between narratives and visual stimuli, recent studies have used text-visual pretrained models to identify the highest relevance scores between image regions and sentences [39] or pattern-matching with fixed references [40, 41]. However, these methods have limitations: local text-image relevance fails to capture global consistency, and pattern-matching is restricted to fixed-reference tasks, making it unsuitable for creative tasks like storytelling. Therefore, tailored and automated approaches are needed for macrostructural analysis of NCD narratives.
This work proposed to investigate specific cognitive and linguistic challenges by analyzing topical shifts, their temporal dynamics, and the coherence of narratives over time. We presented two novel approaches to macrostructural narrative analysis: Dynamic Topic Models (DTM) for temporal analysis of topical evolution; and Text-Image Temporal Alignment Model (TITAN) for cross-modal (i.e., stimulus images and spoken narratives) alignment analysis. This approach aims to reveal cognitive deficits by identifying narrative impairments associated with NCDs, and to provide a deeper understanding of their impact on communication and cognition. The investigations are conducted using the Rabbit Story task (a VSN task) from the CU-MARVEL corpus (a Cantonese corpus), as well as the single-image description task in the ADReSS corpus (an English corpus). Our experiments validated the effectiveness of dynamic topic consistency as a macrostructure metric (F1=0.65, AUC=0.78). The proposed TITAN model achieved the highest performance (F1=0.72, AUC=0.81) in classification tasks, surpassing those of established microstructure and handcrafted macrostructure feature sets. Moreover, experiments of cross-comparison and regression tasks further demonstrated the effectiveness of the proposed dynamic macrostructure modeling approaches for NCD detection.
The paper is organized as follows: Section II reviews related works, Section III introduces corpora used in this study, Section IV introduces the proposed approaches, Section V presents our experimental settings and results, Section VI analyzes and discusses experimental results, and Section VII concludes this work and prospects future works.
II Related Work
Extensive research has been conducted for NCD detection based on spoken language, leveraging various methodologies to identify biomarkers and assess cognitive impairments. This section reviews pertinent literature encompassing both microstructural and macrostructural narrative analysis, collectively laying the groundwork for enhanced narrative analyses. The term “microstructure” refers to fine-grained elements of narrative production, including lexical, syntax, semantics, phonetics, acoustics, and other smaller linguistic units. In contrast, the term “macrostructure” refers to higher-level aspects of narrative construction, such as coherence, cohesion, thematic development, and the overall organization of narratives. Both microstructural and macrostructural analysis are crucial for a comprehensive understanding of cognitive deficits reflected in NCD narratives. However, existing studies often face limitations in capturing the full spectrum of narrative features, particularly the dynamic aspects of narrative construction. By identifying these limitations, this review highlights the gaps the proposed methods aim to fill, specifically by incorporating temporal dynamics and macrostructural consistency, thus advancing the narrative analysis of NCDs.
II-A Microstructural Analysis in NCD Detection
Individuals with NCDs exhibit speech deficits that impact spoken language across both acoustic and linguistic dimensions. Prior research has explored the various aspects of acoustic and linguistic impairments in NCD speech, encompassing prosody (e.g., increased dysfluencies and longer, more frequent pauses), phonetics (e.g., abnormal Mel Frequency Cepstral Coefficients patterns), phonology (e.g., higher rates of mispronunciation of words), lexicon (e.g., reduced type-token-ratio and elevated stopword ratio), syntax (e.g., reduced parse-tree depth), and semantics (e.g., abnormal patterns in word and sentence embeddings) [31, 42, 33, 43, 44, 13, 45, 46, 32, 47, 48, 40, 49, 50, 51]. These microstructural features are essential for understanding the language abilities of individuals with NCDs, providing the foundation for analysis in this domain.
Recent efforts are devoted to identifying highly correlated and significant biomarkers of NCDs from the aforementioned features [52, 53, 54, 55, 56]. For example, Cho et al. [56] linked higher CSF p-tau [^2] CSF p-tau: Cerebrospinal fluid phosphorylated tau, a NCD-related biomarker indicating the burden of neurofibrillary tangle pathology. levels (signifying greater burden of NCD pathology) to specific acoustic and linguistic features: lower preposition and noun ratios, higher pause rates, and shorter speech segment durations. Such microstructural analysis provides valuable insights into local features, but more comprehensive information may be offered by macrostructural approaches that examine broader narrative structure and contexts. Hence, this work extends prior research by examining VSN’s macrostructure such as topic evolution and cross-modal consistency.
(a) The Cookie Theft picture
(b) The Rabbit Story cartoon
Figure 1: Visual stimuli used for NCD detection: (a) Cookie Theft (depicting two kids stealing cookies while a distracted mother is washing dishes, amidst a complex background) and (b) Rabbit Story (showing a girl and her dog searching for a rabbit and encountering various animals and adventures).
II-B Macrostructural Analysis in NCD Detection
While microstructural analysis highlights specific language production impairments associated with NCDs, macrostructural analysis is equally important as it reveals how cognitive deficits affect higher-order cognitive processes like attention, working memory, planning, and organization. This macrostructural perspective is essential for understanding how NCDs affect the construction of coherent, cohesive, and thematically consistent narratives [57, 58, 59, 30].
Macrostructural analysis in the literature generally focuses on features about cohesion (the use of linguistic devices for structural unity) and coherence (logical flow and overall consistency). Individuals with NCDs tend to produce text with wrong co-referencing, inconsistencies in maintaining the main topic, uncompleted or irrelevant information, and disrupted sequencing of events [60, 61, 62, 63, 64].
Traditional macrostructural features were usually extracted manually based on expert experience, such as readability [65], relationship between a proposition and previous one or the global topic [66], or connection between events, macro-propositions and main information units [67].
Recent advances leverage machine learning and large language models (LLMs) for automatic analyze macrostructure of VSNs. Some studies have quantified global semantic coherence by using cosine similarity of consecutive embeddings [37, 45], while others have assessed pragmatic coherence via BLEU and METEOR scores using references [41]. More recent work has adopted LLM as evaluator to automatically predict macro-descriptors [38].
However, as mentioned in Section I, these methods are unsuitable for NCD spoken narratives: similarity scores are confounded by frequent repetitions, reference-based metrics (BLEU, METEOR) are inapplicable due to non-standardized references, and LLM-based evaluation fails to meet interpretability requirements. Additionally, there has been limited investigation into the temporal dynamics of topical evolution and the cross-modality alignment between visual stimuli and speech in VSN tasks. These gaps highlight the need for more advanced approaches to evaluate NCD narratives from both macrostructural and temporal perspectives, addressing the interplay between evolving topics and the integration of visual and spoken elements.
III Corpora
This study is conducted based on the ADReSS (Alzheimer’s Dementia Recognition through Spontaneous Speech) Corpus [68], as well as the CU-MARVEL-RABBIT (Cognitive Assessment Using Machine LeArning EmpoweRed VoicE AnaLysis – RABBIT Story Part) Corpus. These corpora were selected to enable a comprehensive understanding of VSN patterns for NCDs across different language and narrative types. ADReSS collects English recordings from the Cookie Theft picture description task [69], while CU-MARVEL collects Cantonese narratives from the Rabbit Storytelling task [70].
III-A ADReSS
The ADReSS Corpus, developed by the University of Pittsburgh, contains transcripts of spontaneous speech samples from elderly English-speaking adults. The corpus is designed to facilitate research on age-related cognitive decline, focusing on NCDs (specifically, Alzheimer’s disease and mild cognitive impairment). The neurocognitive assessment tasks in this corpus involve the Cookie-Theft Picture Description, as shown in Fig. 1(a), for both Healthy Control (HC) and NCD groups. This task assesses participants’ ability to recognize the elements and produce a coherent narrative based on their spatial relationships and implied events. This corpus consists of audio recordings and transcriptions of interviews with 156 participants in the Cookie-Theft Picture Description task, balanced partitioned into training (54 HCs and 54 NCDs) and testing (24 HCs and 24 NCDs) subsets.
III-B CU-MARVEL-RABBIT
The CU-MARVEL Corpus is designed and collected by our group in Hong Kong, encompassing various cognitive tests. To assess the cognitive capacities of individuals with NCDs in integrating sequential visual information and constructing coherent narratives, we adopt the Rabbit Story task of CU-MARVEL, denoted as “CU-MARVEL-RABBIT” Corpus.
The corpus records Cantonese-speaking participants narrating a story about a girl finding a rabbit through 15 cartoon images, as illustrated in Fig. 1(b). These images generally depict 5 sequential scenes (or topics): (1) the discovery and subsequent loss of the rabbit (images 1-4), (2) an encounter with a squirrel and some ants (images 5-7), (3) a meeting with a goose (images 7-9), (4) a run-in with a cow (9-11), and (5) a reunion with the rabbit followed by a farewell wave to its family (images 12-15).
These images are presented in a booklet, with each image displayed on a separate page. Participants will first review all 15 images to familiarize themselves with the content. Then, they will examine each image individually, providing a description of its content before proceeding to the next image. This process will continue sequentially until the stories for all 15 images have been completed.
Figure 2: Illustrative Cantonese narratives in CU-MARVEL-RABBIT, where English words are roughly translated by GPT-4o for reader’s comprehension only.
The CU-MARVEL-RABBIT Corpus (version “2024-06-13”) consists of 758 subjects, including 467 HCs and 291 individuals with NCDs. Illustrative examples in Fig. 2, where the English translations are coarsely rendered by GPT-4o [71] for reader’s comprehension only, highlight the typical differences between HC and NCD narratives. Specifically, HC tends to capture most visual details (e.g., characters (girl, dog, rabbit, squirrel, goose, cow), objects (tree hollow, rock), and settings (house, garden, forest)) and weaves them into a clear, logical storyline across the images (e.g., rabbit was missing → girl and dog searched in the forest → rabbit was found with its family). In contrast, NCD often misses or misidentifies key details (e.g., squirrel, goose) and exhibits difficulties in establishing coherent connections between events (primarily isolated descriptions).
The data distribution and train-test split are listed in Table I. To benchmark the performance of frontend tools (including speech recognition, word segmentation, and part-of-speech tagging), we extensively annotated a subset of 46 participants (22 HCs, 24 NCDs) from the test set (aligned with [48]), naming “Frontend-Evaluation” set. The CU-MARVEL-RABBIT Corpus records the assessment for each subject by specialists in neurology, using a 0–4 severity scale for cognitive impairment, with lower values corresponding to better cognitive performance. Specifically, values 0-1 indicate HCs and 2, 3, and 4 represent progressively severe NCDs.
TABLE I: Distribution of CU-MARVEL-Rabbit Corpus
Split Gender Label Total
0 1 2 3 4
Train Male 94 107 5 86 12 304
Female 81 145 37 74 17 354
Total 175 252 42 160 29 658
Test Male 10 10 10 10 10 50
Female 10 10 10 10 10 50
Total 20 20 20 20 20 100
IV Approach
In this work, we propose a comprehensive analysis for the automatic detection of Neurocognitive Disorders based on VSNs, encompassing both microstructural and macrostructural analysis. Microstructural analysis, detailed in Section IV-A, focuses on the extraction of acoustic and linguistic features selected from previous literature. On the other hand, macrostructural analysis, presented in Sections IV-B to IV-D, introduces three distinct methodologies to capture global and temporal patterns with varying focal points.
IV-A Traditional Acoustic and Linguistic Analysis
Based on the literature review, the microstructural features have been categorized into two main groups: acoustic features and linguistic features. We select 10 acoustic features and 13 linguistic features from the literature that are basic, universal and suitable for Cantonese in the CU-MARVEL-RABBIT Corpus, as summarized in Table II. Before extracting these features, we preprocess all audios with FullSubNet+ [72] denoiser and peak-normalization to ensure internal consistency and avoid the Clever Hans Effect (where the model exploits irrelevant or unintended shortcuts like noise) [73].
The rows of A1-A10 in Table II list the acoustic features, related to pause (A1-A6), speech duration (A7-A8), articulation rate (A9) and speaking rate (A10). The articulation rate measures phonetic production speed as the number of syllables per voiced duration (excluding pauses), while the speaking rate quantifies overall delivery speed as the number of syllables per total duration (including pauses), thereby reflecting real-time communication fluency [74].
To extract these features, we first employ Voice Activity Detection (VAD) using the pretrained Feedforward Sequential Memory Network, as described in the paper [75], to obtain the timestamps of the voiced segments. With these segment timestamps, a pause is detected by the silence gap between two adjacent voiced segments when the duration exceeds the duration of a syllable for this speaker, which is calculated by dividing the total voiced duration by the total number of syllables for each Cantonese participant.
The linguistic features are extracted from the automatic transcriptions of participants obtained using the pretrained Automatic Speech Recognition (ASR) model SenseVoice-Small [76], which demonstrates a Word Error Rate (WER) of 31.37% on the ADReSS test set and a Character Error Rate (CER) of 34.22% on the Frontend-Evaluation set of the CU-MARVEL-RABBIT corpus. The rows of L1-L13 in Table II show the name and the corresponding description of each feature. To extract these linguistic features from Cantonese transcripts, we utilized pretrained BERT base models for word segmentation [^3] https://huggingface.co/ckiplab/bert-base-chinese-ws and POS auto-tagging [^4] https://huggingface.co/ckiplab/bert-base-chinese-pos, with F1 scores of 82.69% and 85.11% respectively on the Frontend-Evaluation set of the CU-MARVEL-RABBIT corpus.
TABLE II: Microstructural features used in this work.
“%”: “ratio of”, “#”: “number of”, “dur”: “duration”.
ID Feature Name Description
Acoustics
A1 # pauses The number of pauses
A2 Total pause dur all pause dur in seconds
A3 Average pause dur averaged pause dur in seconds
A4 Normalized pause dur Pause dur / articulation rate
A5 Pause frequency # pauses / speech dur
A6 Pause occurrence rate # pauses / # syllables
A7 Total utterance dur all utterance dur in seconds
A8 Average utterance dur averaged utterance dur in seconds
A9 Articulation rate # syllables / voiced dur
A10 Speaking rate # syllables / total dur (include pauses)
Linguistics
L1 # words The number of words
L2 % Stop words # stop words / # words
L3 % Filled pauses # filler words / # words
L4 % Lexical filler # lexical filler / # words
L5 % Backchannel # backchannel words / # words
L6 % Repetition # repeated words / # words
L7 % Adj # adjectives / # words
L8 % Adv # adverbs / # words
L9 % Noun # nouns / # words
L10 % Pronoun # pronouns / # words
L11 % Verb # verb / # words
L12 % Func # functional words / # words
L13 % CTTR # types / rooted(2 * # tokens)
Figure 3: The topic distribution for the HC group (a) and the NCD group (b), and the topic evolution for the HC group (c) and the NCD group (d), as obtained from DTM. (a-b) show the weights of each topic word (top-10) within the topic (from A to E) for the group. (c-d) display average topic proportions with curves and deviations with filled areas. The order of topics A to E is based on the positions of their first peaks in topic evolutions.
IV-B Reference-based Narrative Alignment
A critical part of evaluating the narrative task, particularly within the Rabbit Story context, involves assessing the comprehensiveness of a subject’s narrative. To this end, we identify keywords associated with the task and measure the coverage in the subject’s narratives. Specifically, we rank words from the HCs group (excluding punctuation and stopwords) based on their relevance to visual stimuli. This relevance is quantified through cosine similarity between global image embeddings (from the global special token [CLS]) and text embeddings (via mean pooling), both extracted from the final projection layers of the visual and textual encoders in the Chinese-CLIP model [77], which is pretrained via a contrastive learning paradigm that aligns images with corresponding textual descriptions on diverse Chinese variations (including the Cantonese dialect). As the textual and visual modalities are projected into a shared multimodal embedding space, the Chinese-CLIP model inherently enables cross-modal similarity measurement. These derived top-ranked words are categorized into four semantic groups informed by human prior knowledge of the Rabbit Story, and the Narrative Scoring Scheme (NSS) proposed by [24], including general settings or introduction (e.g., daytime/one day, garden/house/forest), characters or roles (e.g., girl, dog, rabbit), objects (e.g., sofa, box, stone), and actions (e.g., play, find, walk). Coverage (abbreviated as “cvg”, ranging from 0 to 1) for each category is then computed as features for analysis, focusing on the narrative content instead of linguistic form like PoS ratios. For example, the Rabbit Story features 7 characters, but the NCD narrative in Fig. 2(b) only mentions 4 of them, resulting in a “cvg_role” value of 4/7.
While these coverage metrics effectively capture whether local narrative elements are addressed, they lack insights into global coherence and connections across the entire picture set. To address this limitation, metrics commonly used in image captioning tasks, such as BLEU-n (n=1,2,3,4), METEOR, and ROUGE-L, are employed to quantify the degree of alignment between subject narratives and reference descriptions. For this purpose, we engaged six research team members, all thoroughly familiar with the NSS rubric [^5] The NSS rubric evaluates narratives across seven dimensions: introduction, character development, mental and emotional states, referencing, conflict/resolution (or event/reaction), cohesion, and conclusion. Each dimension uses a 0-5 scale, with concrete elaborations and examples provided for the “Proficient (5)”, “Emerging (3)”, and “Minimal/Immature (1)” levels., to craft reference descriptions of the Rabbit Story that exemplify the proficient criteria. The final BLEU-n, METEOR, and ROUGE-L scores are averaged across these references for each subject, providing additional features that reflect global narrative coherence.
In short, we derive two types of features for reference-based analysis: (1) 10 domain-specific coverage metrics, including introduction, character, object, and action words, as well as high-frequency visual-related terms (nouns, verbs, pronouns, adjectives, adverbs, and all types of POS); and (2) 6 expert-aligned narrative similarity scores (BLEU-n, METEOR, ROUGE-L). These reference-based features enable macrostructural analysis by evaluating informativeness and coherence against curated standards. However, as these features derive primarily from word frequencies or local n-gram matching, they are incapable of modeling temporal topic development or the consistency.
IV-C DTM-based Temporal Analysis
Temporal changes or thematic evolution in narratives are critical for understanding the progression and manifestations of NCDs, as these elements reflect higher-level cognitive abilities. To capture these temporal dynamics, we utilize Dynamic Topic Modeling (DTM) [78], which tracks them by explicitly modeling the evolution of topics over time. In contrast to previously mentioned static methods, DTM offers a dynamic perspective on the underlying temporal trends within the narratives.
To understand how DTM captures these temporal dynamics, it is important to first grasp the basics of topic modeling. Topic modeling provides a probabilistic framework to uncover the latent semantic structure within textural data. While the arbitrarily generated text may be randomly distributed across documents, the topic model assumes that words in natural languages are distributed around meaningful semantic topics. Specifically, document generation is assumed to be based on a probabilistic combination of underlying topics, where each topic is represented by a probabilistic distribution over vocabulary words. In the classical Latent Dirichlet Allocation (LDA), these distributions are assumed to be static and drawn from Dirichlet priors.
Dynamic Topic Model (DTM) is an extension of classical topic models. It introduces time dependencies upon the distributions, making it particularly suitable for analyzing temporal changes in narrative content. Formally, DTM models the dynamic nature of topic evolution by introducing time dependency $t$ in topic word distribution $\beta_{k,t}$, where $k$ and $t$ denote the topic and time indices respectively. In this way, it allows the topic distributions to drift over time via a Gaussian process, i.e., $\beta_{t}|\beta_{t-1}\sim\mathcal{N}(\beta_{t-1},\boldsymbol{\Sigma})$, where $\boldsymbol{\Sigma}$ represents the covariance matrix controlling topic transition smoothness.
Then the generative process of DTM proceeds as follows:
Firstly, for each document $d$ in time slice $t$, the model first samples a topic proportion from a Dirichlet prior: $\theta_{d}\sim\text{Dir}(\alpha)$, where $\alpha$ is the Dirichlet concentration parameter controlling the sparsity of topic distributions.
Secondly, for each word position $n$ in document $d$, the model samples a topic assignment from a multinomial distribution parameterized by the document topic proportion: $z_{d,n}\sim\text{Mult}(\theta_{d})$.
Finally, the word is sampled from a multinomial distribution defined by the assigned topic’s word distribution at that specific time slice: $w_{d,n}\sim\text{Mult}(\beta_{t,z_{d,n}})$.
The model parameters could be estimated through variational inference [79] or the Expectation-Maximization [80] algorithms during the training phase.
In order to apply the DTM model to narrative analysis for NCD detection, we evenly split each participant’s ASR transcription into 15 slices, assuming that each image leads to approximately similar number of words in spoken narratives. To mitigate fragmented sentences, we shift the split points to the nearest punctuation boundaries. This segmentation allows us to track the evolution of participants’ narrative topics across the story’s progression, enabling a granular analysis of discourse cohesion by observing how thematic content shifts and develops with each sequential image in the narrative.
We use DTM by inputting the segmented transcription data, where each slice represents a distinct portion of the narrative associated with a time step. The DTM processes this data to identify evolving topics within and across the segmented slices, allowing us to observe how smoothly and coherently topics transition as the story progresses.
Figure 4: (a) The proposed model TITAN. (b) The Attention module of TITAN. (c) The correlation map between image and text embeddings in TITAN, where the 15 textual chunks are manually segmented and aligned to the corresponding images. TITAN takes images and text narratives as inputs to predict NCDs, with Attention using RoPE to focus on important positions.
Fig. 3 illustrates the topic distribution (a, b) and temporal evolution of topic proportions (c, d) for the HC and NCD groups. For the distribution of topics over words, there exists differences between HC and NCD groups, where HC topics over time tend to be more aligned with the order of the images. For example, the topic words of the HC group, shown in Fig. 3(a), include “home” and “garden” in topic “A” (corresponds to settings of images 1-4), “ant”, “tree hollow” and “goose” in topic “B” (corresponds to images 5-8), “cow” and “stone” in topic “C” (corresponds to images 9-10), “fall” and “pond” in topic “D” (corresponds to images 10-11), and “many” (rabbits) and “mother” in topic “E” (corresponds to images 12-15). These topic words generally align with the different scenes along the story. In contrast, the topic words of the NCD group, shown in Fig. 3(b), are more chaotic. For example, “tree hollow” in topic “B” and “ant/squirrel” in topic “C” both predominantly appear in the same scene in images 5 and 6, indicating overlapping topic assignments.
Another interesting finding is that the HC narratives exhibit a cyclical pattern. Specifically, most HC individuals mention “return” and “home” in topic ”E”, even though these elements are not depicted in the final images. This implies that HCs can tie the ending of the story back to its beginning, completing the girl’s quest to find the lost rabbit and bringing the story to a satisfying closure.
For the dynamic proportions of these topics over time, as shown in Fig. 3(c, d), it is observable that the HC group exhibits more distinct peaks (with each topic dominating consecutive slices) and smoother transitions between topics over time (peaks rising sharply after the preceding topic decreases), while the NCD group shows more erratic changes with less pronounced peaks (e.g., reduced peak amplitudes and sporadic dominance).
Based on these findings, we proposed 6 DTM-related statistics for NCD detection. The first metric, termed “topic_consistency”, is derived from the hit rate of the top-10 topic words each time, which are derived from a DTM pretrained on all narratives. This statistic measures how consistently the subject’s topic words align with the top-10 words over time, thus providing insights into the stability and coherence of topics across different groups (HC vs. NCD). For example, if the words in text slice include all top-10 topic words at the same timestamp $t$, then topic_consistency($t$) = 1.0. Higher topic consistency suggests a more stable thematic structure, indicating a potentially healthier cognitive state. The second metric, termed “topic_cycle”, which is simply a boolean value indicating whether the word “home” appears in the latter half of each narrative.
The other four metrics are derived from the dynamic properties of topic evolution, including standard deviation (“topic_variability”), correlation (“topic_temporal_corr”) of the proportions over time, peak-to-peak range (difference between the highest and lowest values, termed as “topic_ptp_range”), and rate of change (averaged Euclidean distance of adjacent topic distributions over time, termed as “topic_change_rate”). These statistics capture topic development of VSNs and could be used as features followed by a downstream classifier for NCD detection, which are interpretable through their well-defined mathematical properties.
IV-D TITAN: Text-Image Temporal Alignment Network
While DTM demonstrates potential in capturing topic evolution patterns, it faces challenges in incorporating cross-modal information in tasks such as picture description and VSNs. This limitation impairs topic representation model, preventing the effective utilization of image information and the handling of cross-modality correspondence. Previous studies have attempted to address these issues by inputting images into multimodal large language models to generate reference descriptions and keywords [41]. However, these approaches often overlook temporal dynamics and fail to capture topic evolution throughout the entire narrative. To overcome these challenges, and drawing inspiration from multimodal metrics like CLIPScore and GROOViST [81, 82] (which measure the alignment between text and images through contrastive learning), we propose the Text-Image Temporal Alignment Network (TITAN). TITAN treats images as dynamic topics and constructs temporal relationships using the Attention mechanism, as illustrated in Fig. 4.
This network comprises several key components:
D1. Feature Extraction
The first step involves extracting features from the input text and images using the pretrained text-visual models, comprising textual-encoder and visual-encoder to process different modalities. In this work, we adopted the visual-encoder ViT and textual-encoder RoBERTa from the pretrained Chinese-CLIP model, as mentioned in Section IV-B. We adopt this model because it can effectively align the embeddings of images and Cantonese text in the semantic space, thus providing insights for narrative coherence (structural dynamics) and thematic consistency (content alignment). Specifically, given an input image $I$ and text $T$, we obtain the embeddings $E_{I}$ (from the [CLS] token) and $E_{T}$ (via mean pooling) from the final projection layer of their respective encoders:
$E_{I}=\text{Chinese-CLIP}_{ViT}(I)\in\mathbb{R}^{J\times H}$
$E_{T}=\text{Chinese-CLIP}_{RoBERTa}(T)\in\mathbb{R}^{K\times H}$
where $H$ is the hidden dimension of the extracted embeddings, $J$ is the number of image frames (e.g., $J=1$ for the ADReSS dataset and $J=15$ for the CU-MARVEL-RABBIT corpus), and $K$ is the temporal length of the investigated text sequence.
D2. Embedding Fusion
The extracted embeddings $E_{I}$ and $E_{T}$ are then concatenated to form a combined representation:
$E_{IT}=[E_{I};E_{T}]\in\mathbb{R}^{(J+K)\times H}$
where [;][;][ ; ] denotes the concatenation operation. This concatenation operation offers at least two advantages. First, it allows the subsequent layers to share the same projection layers for both image and textual embeddings, which is parameter-efficient and feasible since the embeddings are aligned to a common multimodal space during the pre-training stage. Second, inputting the concatenated embeddings into the Attention mechanism enables the model to capture combined information from image-image, image-text, and text-text pairs, thereby enhancing the richness of the contextual features.
D3. Projection Layers
The combined embedding $E_{IT}$ is dimensionally adjusted through multiple projection layers, starting with a linear projection to reduce the embedding size:
$E_{down}=E_{IT}W_{down}^{T}+b_{down}\in\mathbb{R}^{(J+K)\times H^{\prime}}$
where $H^{\prime}$ is the reduced dimension, $W_{down}\in\mathbb{R}^{H^{\prime}\times H}$ and $b_{down}\in\mathbb{R}^{H^{\prime}}$ are the weight matrix and bias of the projection layer, respectively.
After the Attention module, the embedding is then up-projected to its original dimension for residual connections:
$E_{up}=E_{attn}W_{up}^{T}+b_{up}\in\mathbb{R}^{(J+K)\times H}$
where $W_{up}\in\mathbb{R}^{H\times H^{\prime}}$ and $b_{up}\in\mathbb{R}^{H}$ are the parameters of the up-projection layer.
This bottleneck projection structure effectively balances dimensional reduction for computational efficiency and dimensional expansion for capturing complex feature interactions.
D4. Attention Module
As shown in Fig. 4(b), the Attention module aims to capture the temporal dynamics and contextual relationships within and between modalities, by the attending on specific positions of the projected embeddings.
To effectively process dynamic temporal information, we incorporate Rotary Position Embedding (RoPE) [83] with the Attention mechanism. RoPE encodes positional information into token embeddings without disrupting the invariance properties of self-attention, making it well-suited for capturing temporal dependencies. Mathematically, the positional encoding vector for the $i$-th position, $p_{i}\in\mathbb{R}^{2H^{\prime}}$, is defined using sinusoidal functions:
$p_{i}=[\sin(i),\cos(i),\ldots,\sin(\frac{2H^{\prime}}{2}i),\cos(\frac{2H^{\prime}}{2}i)]^{T}$
Each element of $p_{i}$ is derived from sinusoidal functions to capture the position information: $\sin(i)$ and $\cos(i)$ are applied to alternate dimensions, with increasing frequencies across dimensions to capture information at various scales. Given the different temporal strides of images and texts, RoPE is applied separately to each modality with different frequencies, i.e., $p_{i}\in\mathbb{R}^{1,\ldots,H^{\prime}}$ for images and $p_{i}\in\mathbb{R}^{H^{\prime}+1,\ldots,2H^{\prime}}$ for texts.
Given the projected embedding $E_{down}$, we compute the queries, keys, and values as follows:
$Q=E_{down}W_{Q}^{T},K=E_{down}W_{K}^{T},V=E_{down}W_{V}^{T}$
where $W_{Q},W_{K},W_{V}\in\mathbb{R}^{H^{\prime}\times H^{\prime}}$ are learnable parameter matrices for query, key, and value projections respectively. The queries and keys, for image and text parts separately, are then transformed using RoPE to compute position-aware Attention scores. Specifically, for each query $q_{i}$ and key $k_{i}$, the transformed versions are computed as:
$q_{i}^{\prime}=q_{i}\cos(p_{i})+q_{i}^{\perp}\sin(p_{i}),k_{i}^{\prime}=k_{i}\cos(p_{i})+k_{i}^{\perp}\sin(p_{i})$
where $q_{i}^{\perp}$ and $k_{i}^{\perp}$ represent the perpendicular components of the query and key vectors.
Finally, the dot-product attention is computed as follows:
$E_{attn}=\text{softmax}\left(\text{Mask}(Q^{\prime}K^{\prime T})/\sqrt{H^{\prime}}\right)V$
where $Q^{\prime}$ and $K^{\prime}$ are the RoPE-transformed queries and keys. Mask ensures the padding positions of input get near-zero attention weights after the softmax. The scaling factor $\sqrt{H^{\prime}}$ prevents the dot products from growing too large in magnitude. The values remain untransformed to retain the original contents that are being passed through the network.
In VSN tasks such as the Rabbit Story, narrative transcriptions must align with reference images. Participants with NCDs may misidentify animal names (anomia) or miss detailed scenarios (visual perception) in these images. For example, most NCDs misidentify “goose” as “duck”, or skip the event of “goose in the brush” in the seventh image. The attention module in the proposed text-image temporal alignment network is able to perceive potential misalignment and discrepancies cross modalities, and therefore recognize NCD participants with the consideration of their narrative macrostructure.
TABLE III: Classification and Regression Results on the CU-MARVEL-RABBIT corpus. Feature dimensions are shown in parentheses. “*”: the baseline systems; $\downarrow$”: lower is better; “Emb.”: embedding(s). Best results are marked in red, and the second-best are underlined.
Sys. Features Model Classification Regression
[…]
1* Acoustics (10) SVM 0.6000 0.8054 0.4500 0.9000 0.6400 0.1580 0.3244
2* Linguistics (13) SVM 0.5376 0.6596 0.4167 0.7576 0.5700 -0.0051 0.3544
3* Micro (23) SVM 0.5843 0.7571 0.4333 0.8966 0.6300 0.1604 0.3240
4 Reference-based (16) SVM 0.6458 0.7946 0.5167 0.8611 0.6600 0.2112 0.3140
5 DTM-based (6) SVM 0.6522 0.7825 0.5000 0.9375 0.6800 0.1952 0.3172
6 Macro (22) SVM 0.6947 0.8029 0.5500 0.9429 0.7100 0.2233 0.3116
7 All statistics (45) SVM 0.6875 0.7783 0.5500 0.9167 0.7000 0.2367 0.3089
8 Text-Visual Emb. (1024) TITAN 0.7238 0.8120 0.6333 0.8444 0.7100 0.3876 0.3130
D5. Prediction
Finally, the up-projected embedding is temporal pooled with Attention mask ($\in\mathbb{R}^{J+K}$) and passed through a fully connected layer to generate the prediction:
$E_{pool}=\frac{\sum_{i=1}^{J+K}\text{mask}_{i}\cdot E_{up,i}}{\sum_{i=1}^{J+K}\text{mask}_{i}}$
$\hat{y}=\sigma(E_{pool}W_{fc}^{T}+b_{fc})\in\mathbb{R}^{C}$
where $C$ is the prediction dimension (number of classes for classification tasks and 1 for the regression task), $W_{fc}\in\mathbb{R}^{C\times H}$ and $b_{fc}\in\mathbb{R}^{C}$ are the fully connected layer’s weight and bias, and $\sigma$ is an activation function (softmax for classification tasks and identity for the regression task).
Throughout the network, residual connections are employed at each stage to facilitate gradient flow and mitigate the vanishing gradient problem.
The TITAN model, with the combination of feature extraction, embedding fusion, and Attention-based contextual alignment, can effectively integrate text and image information to evaluate the consistency and temporal flow in narratives. As shown in Fig. 4(c), the heatmap illustrates how the model captures the alignment between text chunks (x-axis) and a sequence of images (y-axis). The varying intensity of the colors represents the degree of alignment between the image and textual content, and the diagonal intensity represents the temporal consistency and flow of the narrative. Bright spots indicate high alignment, showing that text chunks and images correlate strongly at those points, while the diagonal progression reflects the sequence in which events or scenes unfold coherently over time. Moreover, compared to Fig. 3(c, d) using discrete, term-based topics (e.g., limited to 5 coarse topics), Fig. 4(c) leverages instance-level image-text matching. This contrast demonstrates that representing evolving topics with images enables a finer-grained temporal alignment between narratives and their corresponding visual concepts. These heatmaps also provide clinicians a clear view into how textual narratives and visual elements are cohesively aligned to support the model’s evaluation of narrative consistency and temporal flow.
V Experiments & Results
V-A Experimental Setup
In this study, we conduct both classification and regression tasks to evaluate the performance of various feature sets and classifier models. Eight systems are designed, each utilizing distinct feature sets. System 1 employs 10 acoustic features, System 2 incorporates 13 linguistic features, and System 3 combines these microstructural features. System 4 utilizes 16 proposed reference-based metrics, and System 5 employs 6 proposed DTM-based features. System 7 combines these macrostructural features, while System 7 integrates all of these microstructural and macrostructural features. As these features are all statistical in nature, we first apply Principal Component Analysis (PCA) with 5 components to reduce dimensionality and then utilize a Support Vector Machine (SVM) for prediction, with a Radial Basis Function (RBF) kernel and soft margin $C=1$. The hyper-parameters “components” of PCA, and “kernel” and “soft margin” of SVM are grid searched using System 7 with all statistical features. System 8, the TITAN model, as depicted in Fig. 4, is connected with a fully-connected linear layer as the task head. To align with PCA in Systems 1 to 7, the projection dimension $H^{\prime}$ of TITAN is also set as 5.
V-B Evaluation Protocol
For the evaluation, we employed multiple metrics across both classification and regression experiments. For the classification task, we used the F1-score and AUC as the primary metrics, supplemented by recall (Rec.), precision (Prec.), and accuracy (Acc.), reported on the positive class (NCD). Despite of AUC, the classification metrics were computed with a default threshold set at 0.5. For regression, we normalized the ground-truth labels to the range [0, 1] and used the coefficient of determination ($R^{2}$, ranging from $\infty$- ∞ to 1) and Root Mean Square Error (RMSE, ranging from 0 to +$\infty$+ ∞) as metrics. To ensure robust evaluation, we averaged the results from the final five epochs of each experiment. Notably, both the ADReSS and CU-MARVEL-RABBIT test sets were unseen and maintained an equal distribution of positive and negative samples, ensuring a balanced evaluation.
V-C Results
Table III presents the results of classification and regression tasks, comparing various feature sets and models using F1 as the primary evaluation metric. Among the systems, the proposed TITAN model (Sys. 8), which utilizes combined text and image embeddings, achieves the best performance in both tasks generally, with an F1 score of 0.7238 and AUC of 0.8120 in classification, and R2 of 0.3876 in the regression task. In contrast, linguistic features-based system (Sys. 2) shows an inferior performance among these systems, which is limited by the ASR, segmentation and tagger systems that are not tailored for colloquial Cantonese narratives.
Another finding is that macrostructural statistics (Sys. 6) outperform microstructural statistics (Sys. 3), achieving F1 scores of 0.6947 versus 0.5843 in classification, and R2 of 0.2233 versus 0.1604 in regression, indicating the efficacy of discourse-level analysis on NCD narratives.
VI Analysis & Discussion
Figure 5: Correlation coefficients between proposed features and NCD labels, highlighting the statistical significance of each feature on the x-axis. The amplitudes of the green (macrostructural) and red (microstructural) bars represent correlation values, indicating that most of the top-ranked features are macrostructural.
VI-A Comparison with previous literature
To evaluate the performance of the models, we conducted a cross-comparison study on the public ADReSS dataset. The results in Table IV demonstrates that the proposed TITAN model (Sys.8) achieves superior performance with the F1 score of 0.8889, exceeding others by at least 3.13% relatively.
TABLE IV: Cross comparison study (binary classification) on ADReSS.
Reference F1 AUC Rec. Prec. Acc.
Yuan’2020 [84] 0.8619 - 0.8540 0.8700 0.8540
Zhu’2023 [39] - - - - 0.8224
Botelho’2024 [38] $0$ 0.8330 0.7690 0.8130
Li’2024 [41] 0.8510 - 0.8330 0.8700 0.8540
Ours (Sys.8) 0.8889 0.9410 0.8333 0.9524 0.8958
VI-B Statistical Features Correlation Ranking
To evaluate the significance of the proposed statistical features, we performed Spearman’s rank correlation test to rank all features based on the absolute values of correlation coefficients. As illustrated in Fig. 5, the correlation distribution indicates that the macrostructural features demonstrate greater importance than microstructural features. Notably, DTM-based metric “topic consistency”, acoustic feature “speaking rate”, and several reference-based coverage metrics ranked relatively high with statistical significance ($p<0.001$), suggesting their potential as indicative biomarkers. Furthermore, many features associated with narrative macrostructural abilities display highly negative correlations with NCDs, whereas most pause-related features exhibit positive correlations with NCDs. These findings underscore the potential utility of these features for effective classification and a deeper understanding of NCD-related characteristics.
VI-C Interpretable Feature Contribution Analysis
To elucidate the role of the proposed features during the model’s decision pathways, we quantify their contribution using SHapley Additive exPlanations (SHAP) values [85]. SHAP is an interpretable method rooted in cooperative game theory that attributes the prediction of a model to each input feature fairly and consistently using Shapley values. Specifically, for each feature $x_{i}\in X$ (the complete feature set), the SHAP value $\phi_{i}$ represents its average differences in the model’s predictions when the target feature is added to all possible subsets $S$ of other features, i.e.
$\phi_{i}=\sum_{S\subseteq X\backslash\{x_{i}\}}w_{i}{[f(S\cup\{x_{i}\})-f(S)]},$
where $w_{i}=|S|!(|X|-|S|-1)!/|X|!$ denotes the weight accounting for subset permutations, ||$\cdot$|| ⋅ | denotes set cardinality, and $f(S)=ln{\frac{P(\hat{y}|S)}{1-P(\hat{y}|S)}}$ is the log-odds ratio of the model’s output probability $P$ based on given inputs $S$.
Figure 6: Global SHAP value distribution (beeswarm plot) across all test samples in CU-MARVEL-RABBIT. Features are ranked vertically by mean absolute SHAP values (most impactful at top). Each point represents a single instance: horizontal position indicates the SHAP value (negative/positive impact), color reflects the standardized feature value (blue = low, red = high), and point density along each feature shows value distribution.
This formulation enables global feature importance analysis through aggregation of SHAP values across the entire unseen test set. Figure 6 presents the global feature contributions ranked by their absolute SHAP values. Notably, the proposed macrostructural features, including “topic_variability”, “topic_change_rate”, “topic_consistency” and “ROUGE-L”, emerged as the top-ranked global contributors. These findings highlight the critical role of macrostructural patterns in distinguishing NCDs from HCs. The consistent directionality of their SHAP values (positive/negative associations with NCD classification) further validates their reliable discriminative power across diverse instances. For example, the NCD narrative in Fig. 2(b) exhibits low topic variability (original value: 0.25, normalized value: -1.71), high topic change rate (original: 0.61, normalized: 0.36), and low topic consistency (original: 0.5, normalized: -1.42), contributing to the model’s prediction of NCD with a high probability (0.99).
VI-D Text-Visual Alignment Differences between NCDs and HCs
Figure 7: Relative differences in text-visual correlation maps between NCD and HC groups: Cool colors represent areas of lower correlation in the NCD group; warm colors represent areas of higher correlation in the NCD group. The text timeline is resampled (via undersampling or interpolation) to match the number of images for visualization.
Fig. 7 presents the relative differences in cross-modal alignments between individuals of HCs and NCDs. To match the varying lengths of text segments with the number of images, the text timeline is resampled (through under-sampling or interpolation) to a fixed length of 15 for visualization. The colormap highlights regions of text-visual correlation divergence, with cool colors indicating diminished alignment in NCD narratives and warm colors representing higher correlation for the NCD group. Notably, the NCD narratives exhibit relatively lower text-visual correlation across the heatmap diagonal, signifying their disrupted cross-modal consistencies in the VSN task. These findings demonstrate TITAN’s ability to model cross-modal consistency for narrative analysis, providing a new dimension for NCD detection.
VI-E Ablation Study of TITAN
We conducted an ablation study to assess the contributions of different components (e.g., “RoPE”, “Project-Up&Down”, “Attention”, and “Image embeddings”) of the proposed TITAN model, as detailed in Table V. By removing each component (denoted as ‘w/o’), we observed their impacts on performance for the NCD detection task. The results show that each component is essential, with the absence of the image embedding (w/o Image emb) causing the most significant performance drop, highlighting its critical role in the model’s efficacy.
TABLE V: Ablation study of TITAN by removing each module (“w/o” means “without”) in the NCD detection task. The best results are marked in red, and the second best are underlined.
Approach F1 AUC Rec. Prec. Acc.
TITAN (Sys. 8) 0.7238 0.8120 0.6333 0.8444 0.7100
w/o RoPE 0.7171 0.8203 0.6167 0.8566 0.7080
w/o Project-Up&Down 0.7218 0.7688 0.6533 0.8066 0.6980
w/o Attention 0.6796 0.7962 0.5833 0.8140 0.6700
w/o Image emb 0.6250 0.7855 0.5000 0.8333 0.6400
VII Conclusion & Future Work
VSN serves as both a production and reflection of integrated cognitive abilities, offering a promising avenue to detect NCDs in an automatic and holistic way. While macrostructural narrative analysis is vital for measuring global coherence in NCDs, the patterns of topic evolution and cross-modal consistency that inherent in VSNs remain underexplored. To address this, we proposed two dynamic macrostructural modeling approaches, DTM-based temporal analysis and the TITAN model, to measure the temporal alignment between the narrative topics and the implicit topics in references or visual stimuli. The DTM-based temporal analysis captures the topic evolution of VSNs, while TITAN further models the referenced dynamic topics from the sequential images in a more fine-grained manner. These approaches were evaluated on the Cantonese CU-MARVEL-RABBIT corpus, with TITAN achieving superior performances in both classification and regression tasks. Additional cross-comparison on the English ADReSS corpus and feature contribution analysis further validated the effectiveness of the dynamic macrostructural modeling approaches.
We anticipate the temporal and macrostructural approaches (like TITAN and DTM-based models) to significantly contribute to the early screening and detection of NCDs, providing a more comprehensive framework for analyzing narratives from these perspectives. While these approaches can further be technically optimized, the concepts underlying them offer new insights into quantifying language use that may indicate cognitive decline, and enhance the understanding of NCDs in VSNs. Unlike approaches that employ detection models as black boxes, our proposed methods offer enhanced interpretability through topic evolution analysis and cross-modality correlation analysis, potentially providing clinicians with more transparent insights into the decision-making process. Moreover, they can be extended to enhance multimodal fusion of more modalities, support multilingual generation due to their language-agnostic nature, and detect other spoken disorders.